OpenStudy (anonymous):

a passenger plane made a trip to las vegas and back. on the trip there it flew 432 mph and on the return trip it went 480mph. how long did the trip there take if the return trip took 9 hours?

OpenStudy (anonymous):

9*480 = 4320 4320/432 = 10 10 Hours

OpenStudy (anonymous):

speed=distance/time 480*9=4320 4320/432=10

hero (hero):

\[\color{green}{\huge{\checkmark}} \]